given 4 red balls and 5 green balls in a box, two balls are taken out without replacement. find the probability of the 2nd ball is red.
Answer #1

P(second ball is red) = P(first ball is red and second ball is red) + P(first ball is green and second ball is red)

= 4/9 x 3/8 + 5/9 x 4/8

= (12 + 20)/(9x8)

= 4/9